# LlamaRestTest: Effective REST API Testing with Small Language Models

### v1

@@ 2.2. REST API Testing and Tools for Enhancing OpenAPI Specifications (OAS)
REST API testing is a critical practice for ensuring the proper functioning of web services (Saleem et al., 2016). Leveraging REST API documents such as OAS, numerous black-box REST API testing tools have emerged (Corradini et al., 2022; Atlidakis et al., 2019; Karlsson et al., 2020; Martin-Lopez et al., 2021; Karlsson et al., 2021; Hatfield-Dodds and Dygalo, 2022; Wu et al., 2022; Kim et al., 2023b; tcases, 2023; Liu et al., 2022). These tools aim to detect issues like internal server errors (e.g., HTTP 500 status code) by systematically exploring API endpoints and their parameters. They primarily rely on the machine-readable parts of API specifications (Kim et al., 2022, 2023a). However, these tools often fall short in complex scenarios, particularly when handling specific data formats or inter-parameter dependencies. For instance, when testing operations in the Ohsome API, they may generate random string values for parameters such as bboxes, bcircles, and bpolys, failing to account for their specific constraints.
@@
Assistant tools like NLP2REST (Kim et al., 2023a) and RESTGPT (Kim et al., 2023c) enhance REST API documents with natural language descriptions, offering better support for REST API testing tools. For example, RESTGPT reads all parameter descriptions in OAS and augments them with machine-readable rules. The key rules include type/format, constraints, example values, and inter-parameter dependencies. Inter-parameter dependencies, as studied by Martin-Lopez et al., refer to the relationships between API parameters where the value of one parameter depends on the value of another, ensuring that valid and contextually appropriate combinations of parameters are generated for accurate API testing (Martin-Lopez et al., 2019). Among these, we particularly focus on example values and inter-parameter dependencies because the type/format and constraints rules are ultimately aimed at generating valid parameters for REST API testing tools.
@@ 2.3. Large Language Models
2.3.1. Overview
Large Language Models (LLMs), such as the Generative Pre-trained Transformer (GPT) series, are at the forefront of advancements in natural language processing (NLP) (Hadi et al., 2023). A language model is essentially a type of artificial intelligence that understands, interprets, and generates text in a manner similar to how humans use language (OpenAI, 2023). These models are trained on extensive collections of written text, enabling them to learn a wide range of linguistic patterns and styles. The GPT series, including well-known models like GPT-3, exemplifies these advanced LLMs (Brown et al., 2020). Trained on diverse and vast datasets, they have demonstrated remarkable ability in producing text that is strikingly human-like, making them valuable for a variety of applications. This has expanded the horizons of NLP, leading to its application in challenging areas ranging from education to customer service (Baidoo-Anu and Ansah, 2023; Debby R. E. Cotton and Shipway, 2024). GPT-3 and its successors have shown the effectiveness of LLMs, particularly in their versatility to handle a wide array of language-related tasks (OpenAI, 2023).
2.3.2. Quantization
Quantization is a technique used to reduce the computational resources required while minimizing the loss in model performance (Jacob et al., 2018). It involves converting the standard floating-point representations of a model’s weights and activations into lower-precision formats, such as 8-bit or 4-bit, or even lower (Dettmers and Zettlemoyer, 2023; Li et al., 2024). The primary advantage of quantization is the significant reduction in memory usage and computational demand (Ali et al., 2024).
2.3.3. Fine-Tuning
Fine-tuning is a technique for adapting pre-trained LLMs for specific tasks, allowing them to perform better on specialized datasets by adjusting their learned representations. A notable advancement in this area is the introduction of Lora (Low-rank Adaptation of Large Language Models), which modifies only a small set of parameters during training, specifically the attention mechanism’s low-rank components. This approach minimizes the computational overhead and reduces the number of parameters updated, facilitating efficient fine-tuning of large models (Hu et al., 2021). Building on the principles of Lora, QLora (Quantized Low-rank Adaptation) incorporates quantization into the low-rank adaptation process to further decrease the memory and computational demands. By reducing the precision of the low-rank factors, QLora maintains performance while enhancing the model’s efficiency, making it ideal for deployment in resource-constrained environments (Dettmers et al., 2024).
@@ 2.4. Motivating Example
REST API operations often present significant challenges for traditional testing tools, particularly when managing complex parameter formats and inter-parameter dependencies. The Ohsome API, which calculates the area of OpenStreetMap (OSM) elements (see Figure 1), illustrates this complexity. Consider the /elements/area operation in the Ohsome API. According to the OpenAPI specification (see Figure 1), this operation requires one of three mutually exclusive boundary parameters: bboxes, bcircles, or bpolys, each with specific and intricate formatting rules. For instance, bboxes requires WGS84 coordinates in the format “lon1,lat1,lon2,lat2”, whereas bcircles requires coordinates along with a radius in meters. The complexity is further compounded by the filter parameter, which requires valid OSM tags, the format parameter, which controls the output format, and the time parameter. Current state-of-the-art testing tools often fail in these scenarios by generating invalid API requests, without considering the parameter value format and the inter-parameter dependencies.
@@
For example, these tools might generate requests that include both bboxes and bcircles, violating the mutual exclusivity requirement, and produce incorrectly formatted coordinates. Although REST API testing assistant tools like RESTGPT enhance the specification by providing more detailed descriptions—such as the correct WGS84 coordinate format—they still struggle with understanding and enforcing inter-parameter dependencies. For instance, RESTGPT fails to recognize the mutual exclusivity of bboxes, bcircles, and bpolys, and it does not provide valid values for the filter parameter, which must adhere to the OSM tagging schema (e.g., “amenity=school” or “building=residential”). These issues reduce test coverage and effectiveness.
@@
LlamaRestTest addresses these challenges with an adaptive two-stage approach involving LlamaREST-EX and LlamaREST-IPD, both fine-tuned for REST API testing purposes. Initially, LlamaRestTest generates inputs randomly. If the request fails, LlamaRestTest uses LlamaREST-IPD and LlamaREST-EX to identify valid parameter sets and values. For example, if the initial request does not specify any boundary parameters, the response may return an error message such as: “The query did not specify any parameter. Please remember: You need to define one of the boundary parameters (bboxes, bcircles, bpolys).” Similarly, if a request contains more than two boundary parameters, the server would respond “The query should not have more than one boundary parameters (bboxes, bcircles, bpolys).
@@
LlamaREST-IPD processes the operation’s parameter descriptions along with the server’s response to identify inter-parameter dependencies. In this case, it detects from the server message that only one boundary parameter can be defined at a time. Using this feedback, LlamaREST-IPD refines the next request, ensuring that only one valid boundary parameter (e.g., bboxes) is included. LlamaREST-EX then enhances this process by generating semantically valid values for the parameters, such as correctly formatted bboxes values (“lon1,lat1,lon2,lat2”) or appropriate OSM tags for the filter parameter. By combining the adaptive learning from LlamaREST-IPD with the input-generation capabilities of LlamaREST-EX, LlamaRestTest can effectively create valid parameter sets that comply with the API’s constraints and formatting rules. In addition to this, we also explore the benefit of efficiency through quantization.
@@ 3. Our Approach
Figure 2 presents an overview of LlamaRestTest. LlamaRestTest is designed to address the challenges of REST API testing, particularly in handling complex inter-parameter dependencies and generating semantically valid inputs. In this section, we discuss the key features of LlamaRestTest, including model selection, database construction, fine-tuning, and quantization, which together enable adaptive, efficient, and effective testing. These components work in synergy to dynamically generate and refine API requests based on server feedback, improving both test coverage and fault detection.
Figure 2. Overview of our approach.
@@ 3.1. Base Model Selection
LlamaRestTest uses language models to identify Inter-Parameter Dependencies (IPDs) and generate example values for API parameters. Assistant tools, such as RESTGPT, which leverages OpenAI’s ChatGPT-3.5 Turbo model to enhance OpenAPI specifications with rules for REST API testing, perform similar tasks, including IPD identification and example value generation. However, the cost—approximately \$10 per specification—poses a significant scalability challenge, especially when considering the additional expenses associated with dynamic server interactions beyond static specifications. Moreover, RESTGPT focuses solely on static specifications, whereas our approach seeks to leverage server response messages dynamically. Given the potential exposure of sensitive information in REST API testing scenarios, we also prioritized using open-source models that can be hosted on user’s computing resource.
Our goal was to employ powerful yet lightweight language models to reduce costs, improve speed, and ensure compatibility with various computing environments, including CPU-based systems. However, we encountered challenges with LLMs, such as Llama3-8B, the smallest in the Llama3 family, which still requires considerable GPU resources (Discussion, 2024). Additionally, Llama3-8b lacked the capability to generate valid IPD rules and example values tailored for REST APIs when tested using prompts from RESTGPT, as shown by our results (Table 3 and Table 4).
To address these issues, we applied fine-tuning to enhance the model’s capabilities for REST API testing and employed quantization techniques to make the model more lightweight. Another reason for selecting Llama3-8B as our baseline model is its architecture’s support for various quantization options, including popular 2-bit, 4-bit, and 8-bit integer formats (Frantar et al., 2022; Yao et al., 2021; Shen et al., 2020). Llama3 is also open-source and offers a GPT-Generated Unified Format (GGUF) transformation option—a binary format designed for fast model loading and saving.
@@ 3.2. Database Construction
Figure 2 illustrates how we constructed the fine-tuning databases. We systematically created two databases: the Inter-Parameter Dependency (IPD) database and the Example (EX) database. The goal of this process was to collect parameter names, descriptions, and associated rules for the IPD database, which includes rules such as Requires, Or, OnlyOne, AllOrNone, ZeroOrOne, and Arithmetic/Relational/Complex inter-parameter dependency rules (Martin-Lopez et al., 2019). For the EX database, we collected parameter names, descriptions, and semantically valid values.
To collect parameters for the IPD database, we sourced inter-parameter dependency data from Martin-Lopez’s IPD database (Martin-Lopez et al., 2019), which systematically catalogs inter-parameter dependencies in REST APIs. To ensure a fair comparison with our benchmarks, we excluded any OpenAPI specifications that were already part of our benchmark dataset. After filtering, we retained 551 parameters with inter-parameter dependencies. These parameters focus primarily on common REST API parameter types, such as query parameters and headers, which typically involve inter-parameter relationships.
For the EX database, we leveraged the APIs-guru dataset (APIs.guru, 2023), which provides a large collection of OpenAPI specifications. We gathered 4,133 OpenAPI specifications from this dataset, excluding those used in our benchmark. These specifications include over 1.8 million parameter sets for REST APIs. We specifically targeted parameters with enumerated values, which are crucial for REST API testing. These parameters have a predefined set of valid options (e.g., json, xml for a format parameter).
To generate IPD rules for the parameters in the IPD database, we utilized the RestGPT framework (Kim et al., 2023c). RestGPT extracts inter-parameter dependency rules from natural language descriptions in the API specifications. After automatic rule generation, we conducted a manual validation process to ensure the generated dependencies accurately reflected the API behavior, including mutually exclusive or dependent parameters.
For the EX database, we focused on extracting parameters with enumerated values. From the API documentation, we systematically extracted the parameter name and the corresponding description as inputs, while the enum values were used as outputs for generating the training dataset. This data was structured such that the parameter name and description served as inputs, while the valid enum values were used as outputs for model fine-tuning.
@@
  • Description: A brief description of the parameter’s purpose, extracted from the OpenAPI specification or generated by RestGPT.
@@
The overall process for constructing the IPD and EX databases is largely automated, ensuring scalability across large datasets. Data extraction, rule generation, and value collection are fully automated for both databases. However, for the IPD dataset, we have incorporated an optional manual validation step. This step allows to have more accurate IPD rules, particularly in cases where inter-parameter relationships are complex. While the system can operate without this manual validation step, incorporating it improves the accuracy of the generated dependency rules and ensures more nuanced cases are handled correctly.
@@ 3.3. Fine-Tuning and Quantization
3.3.1. Training Dataset Preparation
Datasets for IPD and EX were preprocessed and reformatted into a custom tokenized format suitable for fine-tuning. For each parameter, a structured format with special tokens was used to provide context and guide the model’s training. For LlamaREST-IPD, the input sequences follow this structure:
@@ 3.3.1. Training Dataset Preparation
This format includes the parameter’s name and description, followed by the dependency rule in a structured manner, enclosed with special tokens like ‘¡s¿‘ and ‘[INST]‘. Here, ‘¡s¿‘ marks the start of the prompt, ‘[INST]‘ indicates the start of the instruction, ‘[/INST]‘ marks the end of the instruction, and ‘¡/s¿‘ denotes the end of the prompt. The text between ‘¡s¿[INST]‘ and ‘[/INST]‘ is the instruction, which provides the context, and the text after ‘[/INST]‘ presents the expected rule. This setup allows the model to learn both the natural language description and the corresponding logic for inter-parameter dependencies.
@@
This tokenization ensures that the model understands the task of generating example values for each parameter in list format. The input includes the parameter name and description, while the output provides a list of valid example values. We experimented with different tokenization schemes and ultimately chose Llama2 tokens, as they are shorter and more efficient than Llama3 tokens. Since Llama3’s agentic token support was not required for REST API testing tasks, the simpler Llama2 tokens were a better fit. The structured format and custom tokens facilitate efficient parsing of inter-parameter dependencies and example values during training.

### v2

@@ 2.2. REST API Testing and Tools for Enhancing OpenAPI Specifications
REST API testing is a critical practice for ensuring the proper functioning of web services (Saleem et al., 2016). Leveraging REST API documents such as OAS, numerous black-box REST API testing tools have emerged (Corradini et al., 2022; Atlidakis et al., 2019; Karlsson et al., 2020; Martin-Lopez et al., 2021; Karlsson et al., 2021; Hatfield-Dodds and Dygalo, 2022; Wu et al., 2022; Kim et al., 2024a; Kimbrough, 2023; Liu et al., 2022; Stennett et al., 2025; Corradini et al., 2024). These tools aim to detect issues like internal server errors (e.g., HTTP 500 status code) by systematically exploring API endpoints and their parameters. While they primarily rely on machine-readable parts of API specifications (Kim et al., 2022, 2023), these tools often fall short in complex scenarios, particularly when handling specific data formats or inter-parameter dependencies (IPDs). For instance, when testing the Ohsome API, they may generate random string values for parameters like bboxes, bcircles, and bpolys, failing to account for their specific constraints.
@@
To address these limitations, assistant tools such as NLP2REST (Kim et al., 2023) and RESTGPT (Kim et al., 2024b) have emerged to enhance REST API documents by leveraging natural language descriptions. These tools augment OAS with machine-readable rules, including type/format specifications, constraints, example values, and IPDs. IPDs, as studied by Martin-Lopez et al. (Martin-Lopez et al., 2019), are particularly important as they define relationships between API parameters where the value of one parameter affects another, ensuring valid and contextually appropriate parameter combinations for testing. Among these enhancement aspects, we particularly focus on example values and IPDs because the type/format and constraints rules are ultimately aimed at generating valid parameters for REST API testing.
@@ 2.3. Large Language Models
LLMs, such as the Generative Pre-trained Transformer (GPT) series, are at the forefront of advancements in NLP (Hadi et al., 2023). A language model essentially understands, interprets, and generates text in a manner similar to how humans use language (OpenAI, 2023). These models are trained on extensive collections of text, enabling them to learn a wide range of linguistic patterns and styles. The GPT series, including well-known models such as GPT-3, exemplifies these advanced LLMs (Brown et al., 2020). Trained on diverse and vast datasets, they have demonstrated promising ability in producing text that is strikingly human-like, making them valuable for a variety of applications. This has expanded the applicability of NLP to challenging areas ranging from education to customer service (Baidoo-Anu and Ansah, 2023; Debby R. E. Cotton and Shipway, 2024). Recently, LLMs have also been applied for automating software testing tasks, showing promising results in test case generation and bug detection (e.g., (Rao et al., 2023; Lemieux et al., 2023; Pan et al., 2025; Alrashedy and Binjahlan, 2024)).
Quantization is a technique used to reduce the computational resources required while minimizing the loss in model performance (Jacob et al., 2018). It involves converting standard floating-point representations of a model’s weights and activations into lower-precision formats, such as 8-bit, 4-bit, or even lower (Dettmers and Zettlemoyer, 2023; Li et al., 2024). The primary advantage of quantization is the significant reduction in memory usage and computational demand (Ali et al., 2024). This technique is particularly relevant in deploying LLMs, which are known for their massive parameter counts, often reaching hundreds of billions and requiring extensive computational resources. By applying quantization, LLMs can be made more accessible for real-world applications, enabling their deployment on devices with limited resources while maintaining acceptable performance levels. For instance, quantizing a 175B parameter model from 16-bit to 4-bit precision can reduce its memory footprint by up to 75%, making it feasible to be run on consumer-grade hardware (Dettmers and Zettlemoyer, 2023).
Fine-tuning is a technique for adapting pre-trained LLMs for specific tasks, allowing them to perform better on specialized datasets by adjusting their learned representations (Brown et al., 2020). A notable advancement in this area is the introduction of Low-Rank Adaptation (LoRA) techniques, which modify only a small set of model parameters during training, specifically the attention mechanism’s low-rank components. This approach minimizes the computational overhead and reduces the number of parameters updated, facilitating efficient fine-tuning of large models (Hu et al., 2022). Building on the principles of LoRA, Quantized Low-Rank Adaptation (QLoRA) incorporates quantization into the low-rank adaptation process to further decrease the memory and computational demands. By reducing the precision of the low-rank factors, QLoRA maintains performance while enhancing the model’s efficiency, making it ideal for deployment in resource-constrained environments (Dettmers et al., 2023).
@@ 2.4. Motivating Example
REST APIs with complex parameter formats and IPDs present significant testing challenges. The Ohsome API, which calculates the area of OpenStreetMap (OSM) elements, exemplifies such complexity. Consider the /elements/area operation of the API (Figure 1). According to the specification, this operation requires one of three mutually exclusive boundary parameters—bboxes, bcircles, or bpolys—each with specific and intricate formatting rules. For instance, bboxes requires WGS84 coordinates in the format “lon1,lat1,lon2,lat2”, whereas bcircles requires coordinates along with a radius in meters. The API’s complexity stems from multiple parameters: filter (which combines OSM type, geometry type, and tag filters), format (which specifies output as geojson, csv, or json), time (which accepts ISO-8601 timestrings), and various boundary parameters that require specific coordinate formats.
@@
Current state-of-the-art testing tools often generate invalid API requests, without considering the parameter value formats and the IPDs. For example, these tools might generate requests that include both bboxes and bcircles, violating the mutual-exclusivity requirement, and produce incorrectly formatted coordinates. Although REST API testing assistant tools like RESTGPT enhance the specification by providing more detailed descriptions—such as the correct WGS84 coordinate format—they still struggle with understanding and enforcing IPDs. For instance, RESTGPT fails to recognize the mutual exclusivity of bboxes, bcircles, and bpolys, and it does not provide valid values for the filter parameter, which must adhere to the OSM tagging schema (e.g., “amenity=school” or “building=residential”). These issues reduce test coverage and effectiveness.
@@
LlamaRestTest addresses these challenges with an adaptive two-stage approach involving LlamaREST-EX and LlamaREST-IPD, both fine-tuned for REST API testing purposes. Initially, LlamaRestTest generates inputs randomly. If the request fails, LlamaRestTest uses LlamaREST-IPD and LlamaREST-EX to identify valid parameter sets and values. For example, if the initial request does not specify any boundary parameters, the response may return an error message such as: “The query did not specify any parameter. Please remember: You need to define one of the boundary parameters (bboxes, bcircles, bpolys).” Similarly, if a request contains more than two boundary parameters, the server would respond with the error message “The query should not have more than one boundary parameters (bboxes, bcircles, bpolys).
@@
LlamaREST-IPD processes the operation’s parameter descriptions along with the server’s response to identify parameter dependencies. In this case, it detects from the server message that only one boundary parameter can be defined at a time. Using this feedback, LlamaREST-IPD refines the next request, ensuring that only one valid boundary parameter (e.g., bboxes) is included. LlamaREST-EX then enhances this process by generating semantically valid values for the parameters, such as correctly formatted bboxes values (“lon1,lat1,lon2,lat2”) or appropriate OSM tags for the filter parameter. By combining the adaptive learning from LlamaREST-IPD with the input-generation capabilities of LlamaREST-EX, LlamaRestTest can effectively create valid parameter sets that comply with the API’s parameter constraints and formatting rules.
While the effectiveness of LlamaREST-IPD and LlamaREST-EX is important, we also explore the benefit of efficiency through quantization. This is important because LlamaRestTest can perform extensive analysis of server responses—involving LLM inferencing—to refine its test cases adaptively. By applying quantization, we reduce the computational load and inference time, enabling LlamaRestTest to handle these frequent interactions with lower overhead. Our goal is to balance effectiveness (through fine-tuning) and efficiency (via quantization).
[…]
A diagram showing the overview of our approach. Refer to caption
Figure 2. Overview of our approach.
@@ 3. Our Approach
Figure 2 presents an overview of LlamaRestTest. LlamaRestTest is designed to address the challenges of REST API testing, particularly in handling complex IPDs and generating valid inputs. In this section, we discuss the key features of LlamaRestTest, including model selection, database construction, fine-tuning, and quantization, which together enable adaptive, efficient, and effective testing. These components work in synergy to generate and refine API requests dynamically based on server feedback, improving both test coverage and fault detection.
@@ 3.1. Base Model Selection
LlamaRestTest uses LLMs to identify IPDs and generate valid values for API parameters. The closest approach is RESTGPT which leverages OpenAI’s ChatGPT-3.5 Turbo model to enhance OpenAPI specifications with rules for REST API testing, perform similar tasks—i.e., identifying IPDs and generating example values. However, the cost—approximately \$10 per specification—poses a challenge, especially when considering the large number of LLM interactions associated with dynamic server interactions. Moreover, RESTGPT focuses solely on static specifications, whereas our approach seeks to leverage server response messages at runtime. Given also the potential exposure of sensitive information in REST API testing scenarios, we prioritized using open-source models that can be hosted on the user’s computing resource.
Our goal was to employ powerful yet lightweight LLMs to reduce costs, improve speed, and ensure compatibility with various computing environments, including CPU-based systems. However, we encountered challenges with LLMs, such as the Llama3-8B model, which still requires considerable GPU resources (Discussion, 2024). Additionally, Llama3-8b lacked the capability to generate valid IPD rules and example values tailored for REST APIs when tested using prompts from RESTGPT, as shown by our results (Table 3 and Table 4).
To address these issues, we applied fine-tuning to enhance the model’s capabilities for REST API testing and employed quantization techniques to make the model more lightweight. Another reason for selecting Llama3-8B as our baseline model is its architecture’s support for various quantization options, including popular 2-bit, 4-bit, and 8-bit integer formats (Frantar et al., 2023; Yao et al., 2021; Shen et al., 2020). Llama3 is also open-source and offers a GPT-Generated Unified Format (GGUF) transformation option—a binary format designed for fast model loading and saving.
@@ 3.2. Database Construction
Figure 2 illustrates how we constructed the fine-tuning databases. We systematically created two databases: the IPD database and the example (EX) database. The goal of this process was to collect parameter names, descriptions, and associated rules for the IPD database, which includes the Requires, Or, OnlyOne, AllOrNone, ZeroOrOne, and Arithmetic/Relational/Complex IPD rules (Martin-Lopez et al., 2019). For the EX database, we collected parameter names, descriptions, and semantically valid values.
For mining IPD data, we used Martin-Lopez’s IPD database (Martin-Lopez et al., 2019), which contains manually collected dependencies from REST APIs. To ensure a fair evaluation on our benchmark, we excluded any OpenAPI specifications that were included in our benchmark dataset. After filtering, we were left with 551 parameters with IPDs. For creating the EX database, we leveraged APIs-guru (APIs.guru, 2023), which provides a large collection of OAS. We gathered 4,133 OpenAPI specifications from this dataset, excluding those used in our benchmark. These specifications include over 1.8 million parameter sets for REST APIs. We specifically targeted parameters with enumerated values which are predefined set of valid values (e.g., json and xml for a parameter on data format).
To generate IPD rules for the parameters in the IPD database, we used RESTGPT (Kim et al., 2024b). RESTGPT extracts IPD rules from natural language descriptions in API specifications. After extracting IPD rules automatically using RESTGPT, we manually validated the rules. The manual validation involved checking whether the extracted IPD rules accurately represented the intended parameter dependencies based on their descriptions. Out of 551 IPDs, 25 (4.5%) required manual fixing. These were primarily complex cases involving more than two parameters or compound rules. One of the authors, experienced in REST-API testing, performed the validation. The process took approximately five hours in total.
Rules were validated based on their logical consistency with the parameter descriptions. For example, a rule initially extracted as “Requires(A, B) AND Requires(B, C)” might be corrected to “AllOrNone(A, B, C)” after manual validation. Similarly, the rule “OnlyOne(X, Y) OR ZeroOrOne(Z)” might be modified to “OnlyOne(X, Y, Z)” to reflect the intended behavior more accurately. It is important to note that manual validation is a one-time step per API version. The validation step is optional and should have limited effect on the results, especially for our dataset where only 4.5% of IPDs required correction. To support reproducibility, we have shared our models and the validated training dataset in our artifact (Kim, 2023).
For the EX database, we focused on extracting parameters with enumerated values. From the API documentation, we systematically extracted the parameter name, the corresponding description, and the specified enum values to create the training dataset. This data was structured such that the parameter name and description served as inputs, whereas the valid enum values were used as outputs for model fine-tuning.
@@
  • Description: A brief description of the parameter’s purpose, extracted from the OAS or generated by RESTGPT.
@@ 3.3. Fine-Tuning and Quantization
3.3.1. Dataset Preparation
The fine-tuning datasets for IPD detection and example value generation were preprocessed and reformatted into a custom tokenized format suitable for fine-tuning. Initially, we attempted to fine-tune a single model for both tasks. However, our preliminary investigations revealed that this approach led to a poor performance in IPD detection, with the model displaying catastrophic forgetting issues. This led us to the decision to create separate fine-tuned models to improve overall performance.
For each parameter, we used a structured format with special tokens to provide context and guide the model’s training. The LlamaREST-IPD model’s input sequences follow this structure:
@@ 3.3.1. Dataset Preparation
This format includes the parameter’s name and description, followed by the dependency rule in a structured manner, enclosed with special tokens <s> and [INST]. Here, <s> and </s> mark the start and end of the prompt, respectively. Similarly, [INST] and [/INST] indicate the start and end of the instruction. The text between <s>[INST] and [/INST] is the instruction, which provides the context; the text after [/INST] presents the expected rule. This setup allows the model to learn both the natural language description and the corresponding logic for IPDs.
@@
This tokenization ensures that the model understands the task of generating example values for each parameter in list format. The input includes the parameter name and description, while the output provides a list of valid example values. We experimented with different tokenization schemes and ultimately chose Llama2 tokens, as Llama3’s agentic token support is not required for REST API testing tasks, and Llama2 tokens are simpler. The structured format and custom tokens facilitate efficient parsing of IPDs and example values during fine-tuning.